# Decoding-Time Language Model Alignment with Multiple Objectives

### v1

@@ D.1 Definitions
Definition 1 ($f$-divergence [1, 10, 11]).
For probability measures $P$ and $Q$, let $\mu$ be a dominating measure of $P$ and $Q$ (i.e. $P,Q\ll\mu$), and let $p,q$ be the Radon-Nikodym derivative [14] $\frac{dP}{d\mu}$, $\frac{dQ}{d\mu}$ respectively. For simplicity, here we assume $q>0$ almost surely. Then $f$-divergence from $P$ to $Q$ is defined as
@@
Definition 2 (Barrier function [33]).
@@
Definition 3 (Expected calibration error [17, 46]).
@@
Hypothesis 1 (Reduced reward mis-specification [50, 39, 21]).
@@
As the objective is convex and the constraints are affine, we can directly apply the Karush-Kuhn-Tucker conditions [33]:
@@
The lemma is revealed by Theorem 1 in [46]. For completeness, we give a brief proof here. Since $f$ is convex and barrier, we can directly use Lagrange multiplier to solve
@@ D.3 Proof of key theorem
[Eq. 13,14 in [38]]propositionpropviewloss If $I_{f}$ is Reverse KL-divergence, Eq. (2) can be viewed as
@@
This proposition is revealed by Eq. 13,14 in [38]. For completeness, we give a brief proof here. Define $Z(x):=\log\sum_{y\in\mathcal{Y}}\pi_{\operatorname{ref}}(y|x)\exp(\frac{1}{\beta}r(y|x))$, which is a constant. Then we have
@@ Appendix E Implementation details
SFT. For Reddit Summary and Helpful Assistant, we supervisedly fine-tune the Llama2-7B models on the Summarize-from-Feedback dataset, following the practice of [45, 54]; For Safety Alignment, we directly deploy a reproduced model (https://huggingface.co/PKU-Alignment/alpaca-7b-reproduced); For HelpSteer, we supervisedly fine-tune a Mistral-7B model on the HelpSteer dataset, following the practice of [60].
@@
Training hyper-parameters. For PPO, we follow the settings of [54] and train for $100$ batches; for DPO, we follow [60], with PERDEVICE_BATCH_SIZE$=1$ and MAX_LENGTH$=256$.
Inference hyper-parameters. For PPO, we follow the settings of [54] with NUM_BEAMS$=1$; for DPO, we follow [60] with BATCH_SIZE$=4$, MAX_LENGTH$=200$ and NUM_BEAMS$=1$.
@@
Evaluation setups. The evaluation scores are calculated on a down-sampled dataset, by off-shelf reward models. For Reddit Summary and Helpfull Assistant, we uniformly sample a subset of $2$k prompts from the test set, following [54]; for Safety Alignment and HelpSteer, we randomly sample of subset of $200$ prompts from the validation set. The generation configurations are set as identical for all algorithms.
@@ F.1 Motivating example
This motivating experiment is based on Fine-Grained RLHF [51]. We tune two T5-large models $\pi_{1},\pi_{2}$ for relevance and factuality respectively, based on a reproduced SFT model and pre-trained reward models, following the instructions of [51]. And we obtain $\pi_{2}$ via reversing the sign of $Q,K$ matrices of the last two layers of $\pi_{1}$. The preference weightings are set as $w\in\left\{(i/10,1-i/10):i\in\{0,1,\ldots,10\}\right\}$. As Figure 5 shows, though the performance is comparable based on normally trained models, a noticeable lag in the performance of RS emerges after a simple reversal of certain parameters.

### v3

@@ D.1 Definitions
Definition 1 ($f$-divergence fdivergence ; fdivergencecsiszar1 ; fdivergencecsiszar2 ).
For probability measures $P$ and $Q$, let $\mu$ be a dominating measure of $P$ and $Q$ (i.e. $P,Q\ll\mu$), and let $p,q$ be the Radon-Nikodym derivative probtextbook $\frac{dP}{d\mu}$, $\frac{dQ}{d\mu}$ respectively. For simplicity, here we assume $q>0$ almost surely. Then $f$-divergence from $P$ to $Q$ is defined as
@@
Definition 2 (Barrier function convexoptimization ).
@@
Definition 3 (Expected calibration error guo2017calibration ; fDPO ).
@@
Hypothesis 1 (Reduced reward mis-specification modelsoup ; rewardedsoups ; personalizedsoup ).
@@
As the objective is convex and the constraints are affine, we can directly apply the Karush-Kuhn-Tucker conditions convexoptimization :
@@
The lemma is revealed by Theorem 1 in fDPO . For completeness, we give a brief proof here. Since $f$ is convex and barrier, we can directly use Lagrange multiplier to solve
@@ D.3 Proof of key theorem
[Eq. 13,14 in DPO ]propositionpropviewloss If $I_{f}$ is Reverse KL-divergence, Eq. (2) can be viewed as
@@
This proposition is revealed by Eq. 13,14 in DPO . For completeness, we give a brief proof here. Define $Z(x):=\log\sum_{y\in\mathcal{Y}}\pi_{\operatorname{ref}}(y|x)\exp(\frac{1}{\beta}r(y|x))$, which is a constant. Then we have
@@ Appendix E Implementation Details
SFT. For Reddit Summary and Helpful Assistant, we supervisedly fine-tune the Llama2-7B models on the Summarize-from-Feedback dataset, following the practice of trl ; RiC ; For Safety Alignment, we directly deploy a reproduced model (https://huggingface.co/PKU-Alignment/alpaca-7b-reproduced); For HelpSteer, we supervisedly fine-tune a Mistral-7B model on the HelpSteer dataset, following the practice of modpo .
@@
Training hyper-parameters. For PPO, we follow the settings of RiC and train for $100$ batches; for DPO, we follow modpo , with PERDEVICE_BATCH_SIZE$=1$ and MAX_LENGTH$=256$.
Inference hyper-parameters. For PPO, we follow the settings of RiC with NUM_BEAMS$=1$; for DPO, we follow modpo with BATCH_SIZE$=4$, MAX_LENGTH$=200$ and NUM_BEAMS$=1$.
@@
Evaluation setups. The evaluation scores are calculated on a down-sampled dataset, by off-shelf reward models. For Reddit Summary and Helpfull Assistant, we uniformly sample a subset of $2$k prompts from the test set, following RiC ; for Safety Alignment and HelpSteer, we randomly sample of subset of $200$ prompts from the validation set. The generation configurations are set as identical for all algorithms.
@@ F.1 Motivating example
This motivating experiment is based on Fine-Grained RLHF wu2023fine . We tune two T5-large models $\pi_{1},\pi_{2}$ for relevance and factuality respectively, based on a reproduced SFT model and pre-trained reward models, following the instructions of wu2023fine . And we obtain $\pi_{2}$ via reversing the sign of $Q,K$ matrices of the last two layers of $\pi_{1}$. The preference weightings are set as $w\in\left\{(i/10,1-i/10):i\in\{0,1,\ldots,10\}\right\}$. As Figure 5 shows, though the performance is comparable based on normally trained models, a noticeable lag in the performance of RS emerges after a simple reversal of certain parameters.